# Identifying treatment response subgroups in observational time-to-event data

### v4

@@ Abstract
Identifying patient subgroups with different treatment responses is an important task to inform medical recommendations, guidelines, and the design of future clinical trials. Existing approaches for treatment effect estimation primarily rely on Randomised Controlled Trials (RCTs), which are often limited by insufficient power, multiple comparisons, and unbalanced covariates. In addition, RCTs tend to feature more homogeneous patient groups, making them less relevant for uncovering subgroups in the population encountered in real-world clinical practice. Subgroup analyses established for RCTs suffer from significant statistical biases when applied to observational studies, which benefit from larger and more representative populations. Our work introduces a novel, outcome-guided, subgroup analysis strategy for identifying subgroups of treatment response in both RCTs and observational studies alike. It hence positions itself in-between individualised and average treatment effect estimation to uncover patient subgroups with distinct treatment responses, critical for actionable insights that may influence treatment guidelines. In experiments, our approach significantly outperforms the current state-of-the-art method for subgroup analysis in both randomised and observational treatment regimes.
Machine Learning, ICML
@@ 1 Introduction
Randomised controlled trials (RCTs) remain the gold standard for studying heterogeneous treatment effects. By randomly assigning patients to control or treated arms, RCTs eliminate potential confounders when assessing a treatment’s impact on outcomes. However, RCTs are often costly, time-consuming, and restricted to specific patient cohorts, which are likely not to reflect the full spectrum of patients seen in clinical practice (Hernán & Robins, 2016, 2010; Cole & Hernán, 2008). Trial findings may thus not generalise to real-world settings, where treatment choice depends on clinical judgment and patient preferences, among other variables that may remain unobserved.
@@
Figure 1: Subgroup treatment effect discovery in time-to-event observational data. Our work aims to identify subgroups of patients with similar treatment responses to guide clinical practice and design clinical trials. Our method simultaneously models the treatment effect and identifies subgroup while addressing treatment non-randomisation and censoring.
@@
To address these limitations, researchers increasingly leverage large-scale observational datasets which capture diverse, real-world populations. Although observational data offer the potential to identify subgroups that RCTs might miss, they also pose challenges due to non-random treatment assignments and confounding (Benson & Hartz, 2000; Hernán & Robins, 2010; Hernán, 2018). A robust methodological literature in causal inference has emerged to tackle these challenges. Techniques such as inverse probability weighting (Cole & Hernán, 2008), marginal structural models (Robins et al., 2000), and doubly robust estimators (Bang & Robins, 2005) allow for more accurate causal effect estimation under appropriate assumptions.
@@
Prior works in machine learning (ML) have extended these ideas to estimate treatment effects from observational data (Bica et al., 2020; Curth et al., 2021; Louizos et al., 2017). However, most methods focus on (i) average treatment effects at the population level or (ii) individualised treatment effects. These approaches overlook the utility of identifying treatment-effect subgroups —groups with consistently distinct responses to a given intervention. Detecting such subgroups is crucial in translating treatment effect estimates into actionable insights for clinical decision-making and resource allocation.
@@
We address this gap by introducing a novel framework that uncovers patient subgroups with distinct treatment responses using observational time-to-event data. Extending the rich tradition of outcome-guided modelling (Foster et al., 2011) to a flexible, neural network–based architecture. Crucially, our approach does not require explicit parametrisation of time-to-event distributions or treatment effects; it instead leverages a mixture of monotonic neural networks with correction for non-random treatment assignment. In summary, our contributions are:
@@
  • Neural network-based treatment subgroup identification. Our work introduces the first neural network architecture to simultaneously estimate and identify subgroups of treatment effects in observational settings, accommodating both censoring and non-random treatment assignments without parameterising the treatment effects.
  • Novel formalisation. In section 2, we review existing approaches to identify subgroups of treatment effects. Section 3 formalises the problem of subgroup discovery and our strategy for minimising biases arising from non-random treatment assignment.
@@
  • Extensive application on survival time-to-event data. In section 3, we introduce the proposed monotonic neural network implementation, and evaluate its performance on two synthetic datasets in Section 4, an extensive set of alternative settings in App. B and a real-world example described in App. C, benchmarking against several established approaches.
2 Related work
While survival subgrouping has been proposed through mixtures of distributions (Nagpal et al., 2021a, b; Jeanselme et al., 2022) to identify different phenotypes of patients, the ML literature on phenotyping treatment effects with time-to-event outcomes remains sparse. The current focus is on estimating population or individual treatment effects in the static (Shalit et al., 2017; Johansson et al., 2016; Zhang et al., 2020) and survival settings (Curth et al., 2021). While these approaches help to understand the population or individualised treatment effects, they do not identify subgroups that may benefit or be harmed by treatment. Identifying such groups aligns with, and is thus more useful for, drafting medical guidelines that direct treatment to those subgroups most likely to benefit from it.
Tree-based approaches were proposed to address the limitations of step-wise approaches by jointly discovering subgroups and modelling the treatment effect. Instead of traditional splits on the observed outcomes, these causal trees aim to discover homogeneous splits regarding covariates and treatment effects. Su et al. (2009) introduce a recursive population splitting based on the average difference in treatment effect between splits. Athey & Imbens (2015, 2016) improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation. Wager & Athey (2018) agglomerate these causal trees into causal forests for improved ITE estimation. Each obtained split in the decision tree delineates two subgroups of treatment effect (Lipkovich et al., 2011; Loh et al., 2015). Alternatively, McFowland III et al. (2018) propose pattern detection and Wang & Rudin (2022), causal rule set learning to uncover these subgroups. However, all these approaches rely on a local optimisation criterion (Lipkovich & Dmitrienko, 2014) and greedy split exploration. Recently, Nagpal et al. (2020) addressed the local optimisation by constraining the treatment response to a linear form in a mixture of Cox models.
Previous approaches uncover subgroups of treatment effect but consider RCTs with binary outcomes, not the observational setting with survival outcomes that our work explores. At the intersection with survival analysis, Zhang et al. (2017) extend causal trees to survival causal trees, modifying the splitting criterion by measuring the difference in survival estimates between resulting leaves. Similarly, Hu et al. (2021) propose Bayesian additive models and Zhu & Gallego (2020) propose a step-wise approach with propensity weighting to study observational data. Closest to our work, (Jia et al., 2021; Nagpal et al., 2023) propose to uncover subgroups within RCTs with survival outcomes. Jia et al. (2021) propose a mixture of treatment effects characterised by Weibull distributions trained in an expectation-maximisation framework. Similarly, Nagpal et al. (2023) stratify the population into three groups: non-, positive- and negative responders to treatment. An iterative Monte Carlo optimisation is used to uncover these subgroups, characterised by a Cox model with a multiplicative treatment effect. As demonstrated in our work, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational data.
@@ 3.1 Problem setup
Our goal is to uncover a pre-specified $K\in\mathbb{N}$ number of subgroups [^1] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 4. with distinct treatment responses, guided by the observed times of occurrence of an outcome of interest. To this end, our model assigns patients to latent subgroups based on their covariates at the treatment decision time. For each subgroup, we estimate the probability of not observing the event of interest given time under treatment and under control regimes. These two distributions are known as survival functions, and their difference corresponds to the estimated treatment effect for a given group.
Formally, consider the random variables associated with observed covariates $X$, the assigned binary treatment $A$, and the observed event time $T^{\prime}$. Following the potential outcomes formulation, $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ under consistency (Assumption 1) where $T^{\prime}_{1}$ is the potential event time under treatment and $T^{\prime}_{0}$, under the control regime.
Assumption 1 (Consistency).
A patient’s observed event time is the potential event time associated with the observed treatment. Formally, this means $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ where $T^{\prime}$ is the observed event time and $(T^{\prime}_{0},T^{\prime}_{1})$ are the potential event times under the control and treatment regimes respectively.
Central to our problem is the latent, unobserved, subgroup membership $Z$. As formalised in Assumption 2, we assume that the variable $Z$ mitigates the dependence between the covariates $X$ and the observed event time $T^{\prime}$.
Assumption 2 (Mixture mitigation).
@@
This focus differs from the literature interest in estimating Individualised Treatment Effects (ITE), defined as follows (see App. A.1 for derivation from the potential outcomes formalisation):
$\tau(t,x)$ $=\mathbb{P}(T^{\prime}\geq t\mid A=1,X=x)$
$-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$
$=\sum_{k}\tau_{k}(t)\mathbb{P}(Z=k\mid X=x)$ (Under Asm. 2)
@@
The absence of treatment randomisation in observational studies, e.g. clinicians might recommend more aggressive treatment to patients in more severe conditions, results in a covariate shift between the treated and non-treated populations (Curth et al., 2021) as their covariate distributions differ (Bica et al., 2021). This prohibits the estimation of the survival functions through the maximisation of the observed outcomes likelihood.
Under Assumptions 3 and 4, remedies such as re-weighting (Shimodaira, 2000), or penalisation on the dissimilarity between treatment regimes (Johansson et al., 2016; Shi et al., 2019), or a combination of these approaches (Curth et al., 2021; Hassanpour & Greiner, 2019; Hernán & Robins, 2006; Shalit et al., 2017) allows to estimate both factual and counterfactual likelihood from the observed data by penalising the difference between the treated and non-treated populations. In our work, we adopt an Inverse Propensity Weighting (IPW) of the factual likelihood to estimate the overall likelihood, consisting in weighting each observation by the inverse probability of receiving treatment.
Assumption 3 (Ignorability).
@@
Assumption 4 (Overlap / Positivity).
@@
To express the factual likelihood, one needs to tackle a final challenge in time-to-event data: censoring. Patients may not observe the outcome of interest during the study period or exit the study for reasons independent of the outcome of interest. Formally, instead of observing $T^{\prime}$, one observes an indicator identifying whether the event of interest was observed $D$, and the associated observed time of event $T$, with $T:=\min(C,T^{\prime})$ and $D:=\mathbbm{1}(C>T^{\prime})$, where $C$ is the random variable of the (right)-censoring time. When a patient is censored, $T=C$ and $D=0$; otherwise, the event under treatment regime $A$ is observed and $T=T^{\prime}_{A}$ and $D=1$.
@@
Assumption 5 (Non-informative censoring).
The censoring time $C$ is independent of the time of the event of interest $T^{\prime}$ given the covariates $X$. Formally, $T^{\prime}\perp\!\!\!\perp C\mid X$.
Under the previous assumptions, we can maximise the overall likelihood by optimising for the following weighted factual log-likelihood $l_{F}$:
$l_{F}=$ $\sum_{i,d_{i}=1}w_{i}\log\left(\frac{\partial S(t\mid x_{i},a_{i})}{\partial t}\bigg{|}_{t=t_{i}}\right)$
@@
where $i$ is the patient index, $(x_{i},t_{i},a_{i},d_{i})$ are realisations of the associated variables $X,T,A$ and $D$, and $w_{i}$ is the inverse propensity weighting correction for patient $i$. Under Assumption 2, the survival corresponds to the weighted survival over the different subgroups:
@@
$=\sum_{k=1}^{K}\mathbb{P}(Z=k\mid x,a)\mathbb{P}(T^{\prime}\geq t\mid Z=k,a)$ (3)
In summary, Figure 2 describes all variables and dependences assumed in the studied problem with a directed acyclic graph. Given the observed $X$, $A$, $T$ and $D$, our aim is to estimate the latent structure $Z$ and the associated survival distributions under the two potential treatment regimes $T^{\prime}_{0}$ and $T^{\prime}_{1}$.
Figure 2: Graphical representation between covariates ($X$), treatment ($A$) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed.
Figure 3: Causal Survival Clustering (CSC) architecture. Latent parameter $u_{k}$ characterising the subgroup $k$ is inputted in the monotonic network $M$ to estimate the survival under both treatment regimes. $G$ assigns the probability of belonging to each subgroup given the patient’s covariate(s) $x$. To tackle the challenge of treatment assignment bias, the network $W$ estimates the treatment propensity used to weigh the training likelihood. In this context, patient survival is the average of the weighted survival distributions across subgroups under the given treatment.
@@ 3.2 Estimating the quantities of interest
The previous section discussed the quantities one must estimate —here parametrised by neural networks [^2] When interpretability is a key concern, one can consider a linear model for both subgroup assignment and IPW components, as explored in Appendix B.4.— to uncover subgroups of treatment effects: the assignment function $\mathbb{P}(Z\mid X)$, the distributions characterising survival under the two treatment regimes $\mathbb{P}(T^{\prime}\geq t\mid A,Z=k)$, and the IPW weights $w$. Figure 3 illustrates the overall architecture and the neural networks used to estimate these quantities.
@@
Similar to (Jeanselme et al., 2022), a multi-layer perceptron $G$ with a final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
@@
Each subgroup $k$ is represented by a parameter vector $u_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation of the survival functions learnt through backpropagation. The vector $u_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic outcomes [^3] To ensure this constraint, we apply a square function on all weights as proposed in (Jeanselme et al., 2022, 2023). and a final SoftPlus layer to ensure positivity. We train this neural network to model survival through the following transformation — ensuring that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks (Shchur et al., 2020):
$\mathbb{P}(T^{\prime}_{A=a}\geq t\mid Z=k):=\exp(-t\cdot M(u_{k},t)[a])$
@@
Under treatment randomisation, as in RCTs, the previous components would accurately model the survival outcomes under both treatment regimes and identify subgroups of treatment effects by maximising the factual likelihood. As previously discussed, to account for the treatment non-randomisation in observational studies, we weight the factual likelihood using the propensity of receiving treatment estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as:
@@
From this estimated probability, we further truncate the propensity score (Austin & Stuart, 2015) to avoid extreme values for the estimated weights, which would result in unstable treatment effect estimates. In this context, the weights $w_{i}$ are defined as:
@@
$w_{i}:=$ $\max(0.05,\min(w_{i}^{\prime},0.95))$ (4)

### v5

@@ Abstract
Identifying patient subgroups with different treatment responses is an important task to inform medical recommendations, guidelines, and the design of future clinical trials. Existing approaches for treatment effect estimation primarily rely on Randomised Controlled Trials (RCTs), which tend to feature more homogeneous patient groups, making them less relevant for uncovering subgroups in the population encountered in real-world clinical practice. Subgroup analyses established for RCTs suffer from significant statistical biases when applied to observational studies, which benefit from larger and more representative populations. Our work introduces a novel, outcome-guided, subgroup analysis strategy for identifying subgroups of treatment response in both RCTs and observational studies alike. It hence positions itself in-between individualised and average treatment effect estimation to uncover patient subgroups with distinct treatment responses, critical for actionable insights that may influence treatment guidelines. In experiments, our approach significantly outperforms the current state-of-the-art method for subgroup analysis in both randomised and observational treatment regimes.
keywords:
Treatment effect, subgrouping, observational data
Data and Code Availability
Experiments are performed on synthetic data and a publicly available dataset from the Surveillance, Epidemiology, and End Results Program [^1] Available at https://seer.cancer.gov/. The code to reproduce the proposed model and the presented results is available on GitHub [^2] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
Institutional Review Board (IRB)
This research does not require IRB approval as it relies on a publicly available dataset from previously approved studies.
@@ 1 Introduction
Randomised controlled trials (RCTs) remain the gold standard for studying heterogeneous treatment effects. By randomly assigning patients to control or treated arms, RCTs eliminate potential confounders when assessing a treatment’s impact on outcomes. However, RCTs are often costly, time-consuming, and restricted to specific patient cohorts, which are likely not to reflect the full spectrum of patients seen in clinical practice (Hernán and Robins, 2016, 2010; Cole and Hernán, 2008). Trial findings may thus not generalise to real-world settings.
@@
Figure 1: Subgroup treatment effect discovery in time-to-event observational data. Our work aims to identify subgroups of patients with similar treatment responses to guide clinical practice and design clinical trials. Our method simultaneously models the treatment effect and identifies subgroups while addressing treatment non-randomisation and censoring.
@@
To address these limitations, researchers increasingly leverage large-scale observational datasets which capture diverse, real-world populations. Although observational data offer the potential to identify subgroups that RCTs might miss, they also pose challenges due to non-random treatment assignments and confounding (Benson and Hartz, 2000; Hernán and Robins, 2010; Hernán, 2018). A robust methodological literature in causal inference has emerged to tackle these challenges. Techniques such as inverse probability weighting (Cole and Hernán, 2008), marginal structural models (Robins et al., 2000), and doubly robust estimators (Bang and Robins, 2005) allow for more accurate causal effect estimation under appropriate assumptions.
@@
Prior works in machine learning (ML) have extended these ideas to estimate treatment effects from observational data (Bica et al., 2020; Curth et al., 2021; Louizos et al., 2017). However, most methods focus on (i) average treatment effects (ATE), which measure the response at the population level, (ii) conditional average treatment effect (CATE), which reflects average response for an individual given its covariates, or (iii) individualised treatment effects (ITE), which aims to estimate the unobservable counterfactual effect. These approaches overlook the utility of identifying treatment-effect subgroups —groups with distinct responses to a given intervention. Identifying such subgroups is crucial in translating treatment effect estimates into actionable insights for clinical decision-making and resource allocation.
@@
We address this gap by introducing a novel framework that uncovers patient subgroups with distinct treatment responses using observational time-to-event data. Extending the rich tradition of outcome-guided modelling (Foster et al., 2011) to a flexible, neural network–based architecture, our proposed approach addresses the limitations of existing methodology for subgroup discovery (see App. A for a full review). Crucially, our approach does not require explicit parametrisation of time-to-event distributions or treatment effects; it instead leverages a mixture of monotonic neural networks with correction for non-random treatment assignment. In summary, our contributions are:
  • Novel formalisation. In Sec. 2, we formalise the problem of treatment subgroup discovery.
@@
  • Neural network-based treatment subgroup identification. We introduce, in Sec. 2, a neural network architecture to jointly estimate and identify subgroups of treatment effects in observational settings.
@@
  • Extensive application on time-to-event data. We evaluate its performance on synthetic data in Sec. 3, with extensive sensitivity analysis in App. C, and a real-world example described in Sec. 4, benchmarking against several established approaches.
@@ 2.1 Problem setup
Our goal is to uncover a pre-specified $K\in\mathbb{N}$ number of subgroups [^3] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Sec. 3. with distinct treatment responses, guided by the observed times of occurrence of an outcome of interest. To this end, our model assigns patients to latent subgroups based on their covariates at the time of treatment decision. For each subgroup, we estimate the probability of not observing the event of interest over time under treatment and under control regimes. These two distributions are known as survival functions, and their difference corresponds to the estimated treatment effect for a given group.
Formally, consider the random variables associated with observed covariates $X$, the assigned binary treatment $A$, and the observed event time $T^{\prime}$. Following the potential outcomes formulation, $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ under consistency (Asmp. 1) where $T^{\prime}_{1}$ is the potential event time under treatment and $T^{\prime}_{0}$, under the control regime.
Assumption 1 (Consistency)
A patient’s observed event time is the potential event time associated with the observed treatment. Formally, this means $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ where $T^{\prime}$ is the observed event time and $(T^{\prime}_{0},T^{\prime}_{1})$ are the potential event times under the control and treatment regimes, respectively.
Central to our problem is the latent, unobserved, subgroup membership $Z$. As formalised in Asmp. 2, we assume that the variable $Z$ mitigates the dependence between the covariates $X$ and the observed event time $T^{\prime}$.
Assumption 2 (Mixture mitigation)
@@
This focus differs from the literature interest in estimating CATE, connected to the previous quantity as follows (see App. B for derivation):
$\tau(t,x)$ $=\sum_{k}\tau_{k}(t)\mathbb{P}(Z=k\mid X=x)$
@@
The absence of treatment randomisation in observational studies, e.g. clinicians might recommend more aggressive treatment for more severe conditions, results in a covariate shift between the treated and non-treated populations (Curth et al., 2021) as their covariate distributions differ (Bica et al., 2021). This prohibits the estimation of the survival functions through the maximisation of the likelihood of the observed outcomes.
Under the common Asmp. 3 and 4 (Hernán and Robins, 2010; Robins, 1986), methods exist to account for the difference between the treated and non-treated populations. One such method adopted in this work is the Inverse Propensity Weighting (IPW), which consists of weighting each observation by the inverse probability of receiving treatment to estimate the overall likelihood.
Assumption 3 (Conditional ignorability)
@@
Assumption 4 (Overlap / Positivity)
@@
A final challenge exists in time-to-event data: censoring. Patients may not observe the outcome of interest during the study period. Formally, instead of observing $T^{\prime}$, one observes an indicator identifying whether the event of interest was observed $D$, and the associated observed time of event $T$, with $T:=\min(C,T^{\prime})$ and $D:=\mathbbm{1}(C>T^{\prime})$, where $C$ is the random variable of the (right)-censoring time. When a patient is censored, $T=C$ and $D=0$; otherwise, the event under treatment regime $A$ is observed and $T=T^{\prime}_{A}$ and $D=1$.
@@
Assumption 5 (Non-informative censoring)
The censoring time $C$ is independent of the time of the event of interest $T^{\prime}$, given the covariates $X$. Formally, $T^{\prime}\perp\!\!\!\perp C\mid X$.
Under the previous assumptions, common in the causal and survival literature, we can maximise the overall likelihood by optimising for the following weighted factual log-likelihood $l_{F}$:
$l_{F}$ $=\sum_{i,d_{i}=1}w_{i}\log\left(-\frac{\partial S(t\mid x_{i},a_{i})}{\partial t}\bigg|_{t=t_{i}}\right)$
@@
where $i$ is the patient index, $(x_{i},t_{i},a_{i},d_{i})$ are realisations of the associated variables $X,T,A$ and $D$, and $w_{i}$ is the inverse propensity weighting correction for patient $i$. Under Asmp. 2,
@@
$=\sum_{k=1}^{K}\mathbb{P}(Z=k\mid x)\mathbb{P}(T^{\prime}\geq t\mid a,k)$ (3)
Figure 2 summarises all variables and dependencies assumed in the studied problem with a directed acyclic graph. Given the observed $X$, AA, $T$, and $D$, our aim is to estimate the latent structure $Z$ and the associated survival distributions under the two potential treatment regimes $T^{\prime}_{0}$ and $T^{\prime}_{1}$.
Figure 2: Graphical representation between covariates ($X$), treatment (AA) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, AA, $T$ and $D$ are observed.
@@ 2.2 Estimating the quantities of interest
The previous section discussed the quantities one must estimate —here parametrised by neural networks [^4] When interpretability is a key concern, one can consider a linear model for both subgroup assignment and IPW components, as explored in App. C.4.— to uncover subgroups of treatment effects: the assignment function $\mathbb{P}(Z\mid X)$, the distributions characterising survival under the two treatment regimes $\mathbb{P}(T^{\prime}\geq t\mid A,Z=k)$, and the IPW weights $w$. Figure 3 illustrates the overall architecture and the neural networks used to estimate these quantities.
@@
A multi-layer perceptron $G$ with a final Softmax layer assigns a patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
@@
Each subgroup $k$ is represented by a parameter vector $u_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation learnt through backpropagation. The vector $u_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic outcomes as in Jeanselme et al. (2022, 2023) and a final SoftPlus layer to ensure positivity. We train this neural network to model survival through the following transformation — ensuring that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks (Shchur et al., 2020):
$\mathbb{P}(T^{\prime}\geq t\mid A=a,Z=k):=\exp(-t\cdot M(u_{k},t)[a])$
@@
As previously discussed, to account for the treatment non-randomisation in observational studies, we weight the factual likelihood using the propensity of receiving treatment estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as:
@@
From this estimated probability, we further truncate the propensity score (Austin and Stuart, 2015) to avoid extreme values for the estimated weights, which would result in unstable estimates of the treatment effect. In this context, the weights $w_{i}$ are defined as:
@@
$w_{i}:$ $=\max(0.05,\min(w_{i}^{\prime},0.95))$
Figure 3: Causal Survival Clustering (CSC) architecture. Latent parameter $u_{k}$ characterising the subgroup $k$ is inputted into the monotonic network $M$ to estimate the survival under both treatment regimes. $G$ assigns the probability of belonging to each subgroup given the patient’s covariate(s) $x$. The network $W$ estimates the treatment propensity used to account for the treatment assignment bias.